# Approximate predictive control barrier function for discrete-time systems

### v1

@@ 4 MAIN THEORETICAL ANALYSIS
In this section, we provide the main theoretical analysis which provides the foundation of the approximation algorithm in Section 5.1. The contributions are as follows. We first show that any CBF according to Definition 1 implies existence of a CBF according to Definition  4. This result enables analysis of the closed-loop behaviour of the predictive CBF algorithm with respect to possible errors. These errors can be in a first step considered to be exogeneous disturbances, implying inherent robustness, as defined in [14], of the original algorithm. In a second step, these errors will be considered to be errors when approximating the optimal value function $h_{\mathrm{PB}}$, enabling a principled analysis of the impact of regression errors on the closed-loop behaviour of the system.
@@ 4.1 Relation of CBF Definitions
In order to show the existence of a CBF according to Definition 4 given a CBF according to Definition 1, we need to relate the continuous function $h_{\mathrm{PB}}$ and its corresponding decrease $\Delta h_{\mathrm{PB}}$ to class $\mathcal{K}$ comparator functions. For the sake of exposition, we therefore first provide the class $\mathcal{K}$ lower bound and its proof, initially proposed in [13], in a slightly altered manner.
Lemma 1 (reformulated from [13])
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\alpha:[0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$ with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\geq\alpha(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ (11)
Proof 4.3.
@@
$p(r)\!:=\!\begin{cases}0&\text{if }r=0,\\ &\\ \parbox{133.72786pt}{\footnotesize$\min\limits_{y}\;\;\gamma(y)\\ \text{s.t.}\;\;y\in\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$}&\mathrm{otherwise}.\end{cases}$ (12)
Since $\gamma$ is a continuous function and $\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$ is a compact set, the above minimum is attained for all $r\in(0,\overline{r}]$. It holds that
@@
which follows directly from the above minimization as $x$ is a feasible point if $\|x\|_{\mathcal{S}}>0$. Additionally, it holds that $p$ is non-decreasing with increasing $r$, since $\tilde{r}>r$ implies $\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<\tilde{r}\}\subseteq\mathcal{D}\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}<r\}$, i.e., the domain in the constrained minimization problem (second case in (12)) is non-increasing. As $p(r)$ is not necessarily continuous for non-convex sets $\mathcal{S}$, we construct a class $\mathcal{K}$ function $\alpha:[0,\overline{r}]\to\mathbb{R}_{\geq 0}$ which is a lower bound of $p$ on $\mathcal{D}\setminus\mathcal{S}$ similarly to the result in [37, Lemma 2.5]. Consider the partitions $(R_{k+1},R_{k}]$ on the interval $(0,\overline{r}]$ with boundary points $R_{k}:=\overline{r}2^{-k},\mathrm{for}\;k=0,1,2,\dots$ and define $\phi_{k}:=2^{-k}p(R_{k+1})$. Note that $\phi_{k+1}<\phi_{k}$, since $p(r)$ is non-decreasing and $2^{-k-1}<2^{-k}$ $\forall k\geq 0$. In addition, it holds that $\phi_{k}\leq p(r),\forall r\in(R_{k+1},R_{k}]$ as $\phi_{k}=2^{-k}p(R_{k+1})\leq p(R_{k+1})$.
@@
which is piece-wise affine on successive intervals $r\in(R_{k+1},R_{k}]$, continuous, as $\lim_{k\rightarrow\infty}\phi_{k}=0$, strictly increasing and is therefore a class $\mathcal{K}$ function. Furthermore, it holds that $\gamma(r)\leq p(r)$, which implies $\alpha(\|x\|_{\mathcal{S}})\leq\gamma(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ as desired.
@@
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}\rightarrow\mathbb{R}_{\geq 0}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ and $\gamma(x)=0$ for all $\partial_{\mathcal{D}}\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\beta:[0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$, with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\leq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$ (13)
Proof 4.5.
@@
$q(r):=\max\limits_{y}\;\;$ $\gamma(y)$ (14)
s.t. $y\in(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}.$
As $\gamma$ is a continuous function and $(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}$ is a compact set, the above maximum is attained for all $r\in[0,\bar{r}]$. It holds that
$q(\|x\|_{\mathcal{S}})\geq\gamma(x)\;\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$
which follows directly from the fact that $x$ is a feasible point in the above maximisation and the fact that $\gamma(x)=0,\;\forall x\in\partial_{\mathcal{D}}\mathcal{S}$. Additionally, it holds that $q$ is non-decreasing with $r$, since $\tilde{r}>r$ implies $(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>r\}\subseteq(\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S})\setminus\{y\in\mathcal{D}\mid\|y\|_{\mathcal{S}}>\tilde{r}\}$, i.e., the domain in the constrained maximisation problem in (14) is non-decreasing.
A continuous, strictly increasing upper bound for $q(r)$ is then given in the proof of [37, Lemma 2.5] by defining the partitions $[R_{k+1},R_{k})$ with $R_{k}=2^{-k}\bar{r}$ for $k\geq 0$, the quantities $\phi_{k}=q(R_{k})+2^{-k}$ for $k\geq 0$ and $\phi_{-1}=q(\bar{r})+2=\phi_{0}$ and defining the piece-wise affine function
$\beta(r)=\begin{cases}0&\textup{ if }r=0\\ \phi_{k}+\frac{\phi_{k-1}-\phi_{k}}{R_{k}-R_{k+1}}(r-R_{k+1})&\textup{ if }r\in[R_{k+1},R_{k})\\ \phi_{-1}&\textup{ if }r=\bar{r}\end{cases}$ (15)
It then holds that $\beta(0)=0$, $\beta$ is strictly increasing as $\phi_{k-1}\geq\phi_{k}$ and $2^{-k+1}>2^{-k}$ and continuous as $\lim_{r\rightarrow 0^{+}}\beta(r)=0$ as $\lim_{k\rightarrow\infty}\phi_{k}=0$ and $\lim_{r\rightarrow\bar{r}^{-}}\beta(r)=\beta(\bar{r})$, therefore we have $\beta\in\mathcal{K}$.
@@
Figure 1: Illustrative example of the class $\mathcal{K}$ lower and upper bounds in Lemmata 1 and 4.4. The function $\gamma(x)$ is minimised on $\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$ (green) and maximised on $(\mathcal{D}\setminus\textup{int}_{\mathcal{D}}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}$ (grey) to obtain the non-decreasing functions $p(\|x\|_{\mathcal{S}})$ and $q(\|x\|_{\mathcal{S}})$, respectively. Finally, the class $\mathcal{K}$ bounds are constructed via piecewise affine lower and upper bounds.
An illustrative example of the constructed upper and lower bounds is provided in Figure 1. We can now use the results in Lemmata 1 and 4.4, to show that for any function $h$, which is a CBF according to Definition 1, the function $\tilde{h}(x):=\max\{0,h(x)\}$ is a CBF according to Definition 4.
@@
Consider a CBF $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 and a corresponding decrease function $\Delta h$. It holds that the function $\tilde{h}:\mathcal{D}\rightarrow\mathbb{R}_{\geq 0}$ given by $\tilde{h}(x):=\max\{0,h(x)\}$ is a CBF according to Definition 4.
Proof 4.7.
We prove the three statements in Definition 4 for $\tilde{h}(x)=\max\{0,h(x)\}$ based on the fact that $h$ fulfills Definition 1:
(1) From Definition 1, it follows that the safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}=\tilde{\mathcal{S}}=\{x\in\mathbb{R}^{n}\mid\tilde{h}(x)=0\}$ and $\mathcal{D}$ are compact and non-empty.
(2) From Lemma 1, it follows that $\exists\;\alpha_{1}\in\mathcal{K}$, such that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq h(x)=\tilde{h}(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$. As $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)\;\forall x\in\mathcal{D}$. From Lemma 4.4, it then holds that as $h(x)$ is continuous on $\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$ and $h(x)=0\;\forall x\in\partial_{\mathcal{D}}\mathcal{S}$ by definition, $\exists\;\alpha_{2}\in\mathcal{K}$ such that $h(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}_{\mathcal{D}}\mathcal{S}$. With $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\tilde{h}(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}$.
(3) We prove a class $\mathcal{K}$ decrease function for $\tilde{h}$ for all $x\in\mathcal{D}$ by considering the following case distinction, where we denote an optimal input as $u^{*}\in\arg\min_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u)\in\mathcal{D}\}$:
@@
    • $f(x,u^{*})\in\mathcal{D}\setminus\mathcal{S}:$
      $\inf_{u\in\mathcal{U}}\{h(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-h(x){\leq}{-}\Delta h(x)$
      ${\Rightarrow}\!\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\tilde{h}(x){\leq}{-}\Delta h(x)$. As $\Delta h$ fulfils the conditions of Lemma 1, it holds that $\exists\beta\in\mathcal{K}$, such that $\Delta h(x)\geq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ and
      $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\beta_{1}(x)$.
    • $f(x,u^{*})\in\mathcal{S}:$ It holds that $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u)\in\mathcal{D}\}=0<\tilde{h}(x)$. Using the lower bound $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)$ from (2), it holds that $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\tilde{h}(x){\leq}{-}\alpha_{1}(\|x\|_{\mathcal{S}})$
    We can therefore define the function $\alpha_{3}(\|x\|_{\mathcal{S}})=\min\{\beta(\|x\|_{\mathcal{S}}),\alpha_{1}(\|x\|_{\mathcal{S}})\}\in\mathcal{K}$, such that
    $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$
  • $\forall x\in\mathcal{S}:\;\inf_{u\in\mathcal{U}}h(f(x,u))\leq 0$
    $\Rightarrow\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u)){\mid}f(x,u){\in}\mathcal{D}\}{-}\underbrace{\tilde{h}(x)}_{=0}{\leq}\underbrace{{-}\alpha_{3}(\|x\|_{\mathcal{S}})}_{=0}$ where we used $h(x)\leq\tilde{h}(x)=0\;\forall x\in\mathcal{S}$.
@@
$\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u))\mid f(x,u){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$
@@
4.2 Inherent robustness
In this section, we use the relation in Theorem 4.6 to analyse the inherent robustness properties of the PCBF algorithm with respect to external disturbances. With inherent robustness, we denote that a controller, which was designed with nominal system dynamics (1) maintains robust stability guarantees when the closed-loop system is subject to disturbances. Consider therefore the following difference inclusion
$x(k+1)$ $\in F(x(k),w(k))$ (16)
=$=\{x^{+}\in\mathbb{R}^{n}\mid$ $x^{+}=\tilde{f}(x(k),u,w(k))$
$u\in\mathcal{U}_{h_{\mathrm{PB}}}\},$
where $\tilde{f}:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\rightarrow\mathbb{R}^{n}$ is continuous with $\tilde{f}(x,u,0)=f(x,u)$, $w(k)\in\mathbb{R}^{p}$ is an exogeneous disturbance and $\mathcal{U}_{h_{\mathrm{PB}}}$ is the set of safe inputs induced by $h_{\mathrm{PB}}$ according to Definition 2. The aim is to show that if in closed-loop, the inputs $u(k)$ are selected from the set $\mathcal{U}_{h_{\mathrm{PB}}}$, the closed-loop system is input-to-state stable (ISS) with respect to the disturbance $w$. For input-to-state stability, we use the following comparator function definition.

### v2

@@ 5 MAIN THEORETICAL ANALYSIS
In this section, we provide the main theoretical analysis that provides the foundation of the approximation algorithm in Section 4.1. We first show that every CBF according to Definition 1 implies existence of a CBF according to Definition 4. This result enables analysis of the closed-loop behaviour of the predictive CBF algorithm in the presence of approximation errors as well as exogenous disturbances.
@@ 5.1 Relation of CBF Definitions
In order to show the existence of a CBF according to Definition 4 given a CBF according to Definition 1, we need to relate the continuous function $h_{\mathrm{PB}}$ and its corresponding decrease $\Delta h_{\mathrm{PB}}$ to class $\mathcal{K}$ comparison functions. We first recall the class $\mathcal{K}$ lower bound proposed in [13].
Lemma 5.6 (reformulated from [13]).
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\mathcal{S}\rightarrow\mathbb{R}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\alpha:[0,\bar{r}]\rightarrow\mathbb{R}$ with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\geq\alpha(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}.$ (16)
Proof 5.7.
@@
$\hskip-2.84544ptp(r)\!:=\!\begin{cases}0&\hskip-14.22636pt\text{if }r=0,\\ &\\ \parbox{133.72786pt}{\footnotesize$\min\limits_{x}\;\;\gamma(x)\\ \text{s.t.}\;\;x\in\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$}&\hskip-14.22636pt\mathrm{otherwise}.\end{cases}$ (17)
Since $\gamma$ is a continuous function and $\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$ is a compact set, the above minimum is attained for all $r\in(0,\overline{r}]$. It holds that
@@
which follows directly from the above minimization as $x$ is a feasible point if $\|x\|_{\mathcal{S}}>0$. Additionally, it holds that $p$ is non-decreasing with increasing $r$, since $\tilde{r}>r$ implies $\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<\tilde{r}\}\subseteq\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$, i.e., the domain in the constrained minimization problem (second case in (17)) is non-increasing. As $p(r)$ is not necessarily continuous for non-convex sets $\mathcal{S}$, we construct a class $\mathcal{K}$ function $\alpha:[0,\overline{r}]\to\mathbb{R}_{\geq 0}$ which is a lower bound of $p$ on $\mathcal{D}\setminus\mathcal{S}$ similarly to the result in [37, Lemma 2.5]. Consider the partitions $(R_{k+1},R_{k}]$ on the interval $(0,\overline{r}]$ with boundary points $R_{k}:=\overline{r}2^{-k},\mathrm{for}\;k=0,1,2,\dots$ and define $\phi_{k}:=2^{-k}p(R_{k+1})$. Note that $\phi_{k+1}<\phi_{k}$, since $p(r)$ is non-decreasing and $2^{-k-1}<2^{-k}$ $\forall k\geq 0$. In addition, it holds that $\phi_{k}\leq p(r),\forall r\in(R_{k+1},R_{k}]$ as $\phi_{k}=2^{-k}p(R_{k+1})\leq p(R_{k+1})$.
@@
which is piece-wise affine on successive intervals $r\in(R_{k+1},R_{k}]$, continuous, as $\lim_{k\rightarrow\infty}\phi_{k}=0$, strictly increasing and is therefore a class $\mathcal{K}$ function. Furthermore, it holds that $p(\|x\|_{\mathcal{S}})\leq\gamma(x)$, which implies $\alpha(\|x\|_{\mathcal{S}})\leq\gamma(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ as desired.
@@
Consider the compact and non-empty sets $\mathcal{S}$ and $\mathcal{D}$, with $\mathcal{S}\subset\mathcal{D}$ and a continuous function $\gamma:\mathcal{D}\setminus\operatorname{int}\mathcal{S}\rightarrow\mathbb{R}$, with $\gamma(x)>0$ for all $x\in\mathcal{D}\setminus\mathcal{S}$ and $\gamma(x)=0$ for all $x\in\partial\mathcal{S}$. Then there exists a $\mathcal{K}$-function $\beta:[0,\bar{r}]\rightarrow\mathbb{R}_{\geq 0}$, with $\bar{r}:=\max_{x\in\mathcal{D}}\|x\|_{\mathcal{S}}$ such that
$\gamma(x)\leq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}\mathcal{S}.$ (18)
Proof 5.9.
@@
$q(r):=\max\limits_{x}\;\;$ $\gamma(x)$ (19)
s.t. $x\in(\mathcal{D}\setminus\operatorname{int}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}.$
As $\gamma$ is a continuous function and $(\mathcal{D}\setminus\operatorname{int}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}$ is a compact set, the above maximum is attained for all $r\in[0,\bar{r}]$. It holds that
$q(\|x\|_{\mathcal{S}})\geq\gamma(x)\;\;\forall x\in\mathcal{D}\setminus\operatorname{int}\mathcal{S}$
which follows directly from the fact that $x$ is a feasible point in the above maximisation and the fact that $\gamma(x)=0,\;\forall x\in\partial\mathcal{S}$. Additionally, it holds that $q$ is non-decreasing with $r$, since $\tilde{r}>r$ implies $(\mathcal{D}\setminus\operatorname{int}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}\subseteq(\mathcal{D}\setminus\operatorname{int}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>\tilde{r}\}$, i.e., the domain in the constrained maximisation problem in (19) is non-decreasing.
A continuous, strictly increasing upper bound for $q(r)$ is then given in the proof of [37, Lemma 2.5] by defining the partitions $R_{k+1},R_{k})$ with $R_{k}=2^{-k}\bar{r}$ for $k\geq 0$, the quantities $\phi_{k}=q(R_{k-1})+2^{-k}$ for $k\geq 0$ with $R_{-1}=\bar{r}$ and the piece-wise affine function
$\beta(r)=\begin{cases}0&\textup{ if }r=0\\ \phi_{k+1}+\frac{\phi_{k}-\phi_{k+1}}{R_{k}-R_{k+1}}(r-R_{k+1})&\textup{ if }r\in[R_{k+1},R_{k})\\ \phi_{0}&\textup{ if }r=\bar{r}.\end{cases}$
It then holds that $\beta(0)=0$, $\beta$ is strictly increasing as $\phi_{k}\geq\phi_{k+1}$ as $2^{-k}>2^{-k-1}$ and $q$ is increasing, and $\beta$ is continuous as $\lim_{r\rightarrow 0^{+}}\beta(r)=0$ as $\lim_{k\rightarrow\infty}\phi_{k}=0$ and $\lim_{r\rightarrow\bar{r}^{-}}\beta(r)=\beta(\bar{r})$, therefore we have $\beta\in\mathcal{K}$.
@@
Figure 1: Illustrative example of the class $\mathcal{K}$ lower and upper bounds in Lemmata 5.6 and 5.8. The function $\gamma(x)$ is minimised on $\mathcal{D}\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}<r\}$ (green) and maximised on $(\mathcal{D}\setminus\textup{int}_{\mathcal{D}}\mathcal{S})\setminus\{x\in\mathcal{D}\mid\|x\|_{\mathcal{S}}>r\}$ (grey) to obtain the non-decreasing functions $p(\|x\|_{\mathcal{S}})$ and $q(\|x\|_{\mathcal{S}})$, respectively. Finally, the class $\mathcal{K}$ bounds are constructed via piecewise affine lower and upper bounds.
An illustrative example of the constructed upper and lower bounds is provided in Figure 1. We can now use the results in Lemmata 5.6 and 5.8, to establish a relation between CBFs in Definitions 1 and 4.
@@
Consider a CBF $h:\mathcal{D}\rightarrow\mathbb{R}$ according to Definition 1 and a corresponding decrease function $\Delta h$. It holds that the function $\tilde{h}(x):=\max\{0,h(x)\}$, with $\tilde{h}:\mathcal{D}\rightarrow\mathbb{R}$, is a CBF according to Definition 4.
Proof 5.11.
We prove the three statements in Definition 4 for $\tilde{h}(x)=\max\{0,h(x)\}$ based on the fact that $h$ satisfies Definition 1:
1) From Definition 1, it follows that the safe set $\mathcal{S}=\{x\in\mathbb{R}^{n}\mid h(x)\leq 0\}=\tilde{\mathcal{S}}=\{x\in\mathbb{R}^{n}\mid\tilde{h}(x)\leq 0\}$ and $\mathcal{D}$ are compact and non-empty.
2) From Lemma 5.6, it follows that $\exists\;\alpha_{1}\in\mathcal{K}$, such that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq h(x)=\tilde{h}(x)\;\forall x\in\mathcal{D}\setminus\mathcal{S}$. As $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)\;\forall x\in\mathcal{D}$. From Lemma 5.8, as it holds that as $h(x)$ is continuous on $\mathcal{D}\setminus\operatorname{int}\mathcal{S}$ and $h(x)=0\;\forall x\in\partial\mathcal{S}$ by definition, $\exists\;\alpha_{2}\in\mathcal{K}$ such that $h(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\operatorname{int}\mathcal{S}$. With $\tilde{h}(x)=0\;\forall x\in\mathcal{S}$, it holds that $\tilde{h}(x)\leq\alpha_{2}(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}$.
3) We prove a class $\mathcal{K}$ decrease for $\tilde{h}$ for all $x\in\mathcal{D}$ by considering the following case distinction, denoting an optimal input as $u^{*}\in\arg\min_{u\in\mathcal{U}}\{h(f(x,u,0))\mid f(x,u,0)\in\mathcal{D}\}$:
@@
    • $f(x,u^{*},0)\in\mathcal{D}\setminus\mathcal{S}:$
      $\inf_{u\in\mathcal{U}}\{h(f(x,u,0)){\mid}f(x,u,0){\in}\mathcal{D}\}{-}h(x){\leq}{-}\Delta h(x)$
      ${\Rightarrow}\!\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0)){\mid}f(x,u,0){\in}\mathcal{D}\}{\text{-}}\tilde{h}(x){\leq}{-}\Delta h(x)$.
    • As $\Delta h$ fulfils the conditions of Lemma 5.6, it holds that
      $\exists\beta\in\mathcal{K}$, such that $\Delta h(x)\geq\beta(\|x\|_{\mathcal{S}})\;\forall x\in\mathcal{D}\setminus\mathcal{S}$ and
      $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0)){\mid}f(x,u,0){\in}\mathcal{D}\}\text{-}\tilde{h}(x){\leq}{-}\beta(\|x\|_{\mathcal{S}})$.
    • $f(x,u^{*},0)\in\mathcal{S}:$
      It holds that $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0))\mid f(x,u,0)\in\mathcal{D}\}{=}0$.
      Using the lower bound $\alpha_{1}(\|x\|_{\mathcal{S}})\leq\tilde{h}(x)$ from step 2),
      it holds that
      $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0)){\mid}f(x,u,0){\in}\mathcal{D}\}{-}\tilde{h}(x){\leq}{-}\alpha_{1}(\|x\|_{\mathcal{S}})$
    We can therefore define the function $\alpha_{3}\in\mathcal{K}$ with
    $\alpha_{3}(\|x\|_{\mathcal{S}})=\min\{\beta(\|x\|_{\mathcal{S}}),\alpha_{1}(\|x\|_{\mathcal{S}})\}$, such that
    $\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0))\mid f(x,u,0)\in\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}})$
  • $\forall x\in\mathcal{S}:\;\inf_{u\in\mathcal{U}}h(f(x,u,0))\leq 0$
    $\Rightarrow\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0)){\mid}f(x,u,0){\in}\mathcal{D}\}{-}\underbrace{\tilde{h}(x)}_{=0}{\leq}\underbrace{{-}\alpha_{3}(\|x\|_{\mathcal{S}})}_{=0}$
    where we used $h(x)\leq\tilde{h}(x)=0\;\forall x\in\mathcal{S}$.
@@
$\inf_{u\in\mathcal{U}}\{\tilde{h}(f(x,u,0))\mid f(x,u,0){\in}\mathcal{D}\}-\tilde{h}(x)\leq-\alpha_{3}(\|x\|_{\mathcal{S}}).$
@@
5.2 ISS of the Approximate PCBF Algorithm
We can now establish theoretical guarantees of the closed-loop system under the proposed algorithm despite possible approximation errors and exogenous disturbances. First, we show that the optimisation problem (11) is feasible using the maximum decrease (12) for any $x\in\mathbb{R}^{n}$. Thereby a control input can always be computed using the provided algorithm, even if the state lies outside the theoretical region of attraction $\mathcal{D}_{\mathrm{PB}}$, which may be a subset of the true region of attraction of the PCBF algorithm in [12].
Proposition 5.14.
The optimisation problem (11), where $\Delta\hat{h}_{\mathrm{PB}}(x)$ is defined in (12) is feasible for any $x(k)\in\mathbb{R}^{n}$ and $c_{\Delta}\in(0,1]$ if $\Delta\hat{h}_{\mathrm{PB}}(x)\geq 0$ any $c_{\Delta}=1$ otherwise.
Proof 5.15.
The optimisation problem (12) is feasible for any $x\in\mathbb{R}^{n}$ by choosing any $u\in\mathcal{U}$. Given an optimal solution $u^{*}$ to (12), it holds that $u^{*}$ is feasible in (11) as $u^{*}\in\mathcal{U}$ and for any $x\in\mathbb{R}^{n}$, $\hat{h}_{\mathrm{PB}}(f(x,u^{*},0))-\hat{h}_{\mathrm{PB}}(x)=-\max_{u\in\mathcal{U}}(\hat{h}_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(f(x,u,0)))\leq-c_{\Delta}\Delta\hat{h}_{\mathrm{PB}}(x)$ for all $c_{\Delta}\in(0,1]$ when $\Delta\hat{h}_{\mathrm{PB}}(x)\geq 0$ and for $c_{\Delta}=1$ if $\Delta\hat{h}_{\mathrm{PB}}(x)<0$.
Next, we can show that the closed-loop system is ISS with respect to the approximation errors of $\hat{h}_{\mathrm{PB}}$ in $\mathcal{D}_{\mathrm{PB}}$ if the approximation error is bounded.
Assumption 3
This assumption holds for any continuous regressor $\hat{h}_{\mathrm{PB}}$, such as neural networks with continuous activation functions. Due to continuity of $\hat{h}_{\mathrm{PB}}$ and $h_{\mathrm{PB}}$ as well as compactness of $\mathcal{D}_{\mathrm{PB}}$, a maximum $\epsilon_{h}$ exists for $\max_{x\in\mathcal{D}_{\mathrm{PB}}}|h_{\mathrm{PB}}(x)-\hat{h}_{\mathrm{PB}}(x)|$ according to the extreme value theorem [8, Proposition A.7]. In order to analyse robust stability properties of the closed-loop, we require the following definition of robust positive invariance.
Definition 5.16 (Definition 2.2 in [52]).
A set $\mathcal{A}\subset\mathbb{R}^{n}$ is robustly positively invariant for a system $x(k+1)\in F(x(k),w(k))$ if it holds for all $w(k)\in\mathcal{W}$, for some compact set $\mathcal{W}$ containing all possible disturbance realisations, that
$x(k)\in\mathcal{A}\Rightarrow F(x(k),w(k))\subseteq\mathcal{A}.$
For ISS, we use the following comparison function definition.